OpenStudy (anonymous):

If f(x) = (1)/(sqr(x-1)) Find: (f(x) - f(2))/(x-2)

OpenStudy (anonymous):

=1/sqrt(x-1)-1/(x-2)

OpenStudy (anonymous):

\[((\sqrt{x-1}) - x - 1) / (x - 2)(x - 1)\]

OpenStudy (anonymous):

That's what I came out with. How did you get your answer?

OpenStudy (anonymous):

f(2)=1

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

(1/sqrt(x-1)-1)/(x-2)

OpenStudy (anonymous):

@AJ01 you can simplify even more

OpenStudy (anonymous):

OK, that's what I got so far

OpenStudy (anonymous):

(1-sqrt(x-1))/(x-2)(sqrt(x-1)

OpenStudy (anonymous):

My calculus book says: (-1)/(sqr(x-1))(1+(sqr(x-1))

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ \sqrt{x-1} }-1 }{ x-1 }\]

OpenStudy (anonymous):

Denominator is x - 2

OpenStudy (anonymous):

@Mimi_x3 dude come here

OpenStudy (anonymous):

ok sorry about that..... comment denominator

OpenStudy (anonymous):

srqt(x-1)

OpenStudy (mimi_x3):